(x^2+y^2)y prime =y^2


1
Expert's answer
2021-02-24T07:01:24-0500

Given: "(x^2+y^2)y'=y^2"

"\Rightarrow (x^2+y^2)\frac{dy}{dx}=y^2"

"\Rightarrow \frac{dx}{dy}=\frac{x^2+y^2}{y^2}=\frac{x^2}{y^2}+1"

Take the substitution "x=vy\Rightarrow \frac{dx}{dy}=v+y\frac{dv}{dy}"

Then the given differential equation becomes

"v+y\frac{dv}{dy}=\frac{v^2y^2}{y^2}+1=v^2+1"

"\Rightarrow y\frac{dv}{dy}=\frac{v^2y^2}{y^2}+1=v^2-v+1"

Using separation of variables, we get

"\frac{dv}{v^2-v+1}=\frac{dy}{y}"

Integrating on both sides, we get

"\int \frac{dv}{v^2-v+1}=\int \frac{dy}{y}"

"\int \frac{dv}{v^2-v+\frac{1}{4}-\frac{1}{4}+1}=\int \frac{dy}{y}"

"\int \frac{dv}{(v-\frac{1}{2})^2+(\frac{\sqrt{3}}{2})^2}=\int \frac{dy}{y}"

"\frac{1}{\sqrt{3}/2}arctan(\frac{v-\frac{1}{2}}{\sqrt{3}/2})=ln(y)+c"

"\frac{2}{\sqrt{3}}arctan(\frac{2v-1}{\sqrt{3}})=ln(y)+c"

Substituting "v=\frac{x}{y}" , we get

"\frac{2}{\sqrt{3}}arctan(\frac{2(x/y)-1}{\sqrt{3}})=ln(y)+c"